# Novel Setup for Detecting Short-Range Anisotropic Corrections to Gravity

### v2

@@ Abstract
I Introduction
@@ II Proposed Experiment
Figure 2(b), on the other hand, shows an attractor with $N$ fins and a pendulum with $N-1$ fins. In this case, to within a first order approximation, the net torque on the pendulum due to the isotropic component of the gravitational force vanishes for a large number of attractor orientations. As discussed below, the suppression of the isotropic component of the gravitational torque is enhanced as $N$ is increased. Note that the angular extent of the fins differ in Figs. 2(a) and (b). In (a), the fins occupy half of the available area and in (b) they occupy $75\%$ of the available area. The reason for using a larger fin size with the differential pendulum is discussed in Sec. IV.
@@ III Isotropic Gravity
The calculated torque has the expected periodicity of $2\pi/N$ and is consistent with the an independent calculation of the net torque using the numerical methods described in Ref. [41].
@@ IV Anisotropic Gravity
For the isotropic torque calculation presented in Sec. III, the pendulum and attractor fins each occupied half of the total available area $\pi(R^{2}-r_{0}^{2})$. Figure 2(a) shows a top view of this geometry for the $N=4$ case. For the case of the differential torsion pendulum with $N-1$ pendulum fins and $N$ attractor fins, we found that the desired anisotropic torque vanished when the fins occupied half of the available area. The Supplemental Material presents a systematic study of the magnitude of the anisotropic torque as a function of the angular size of the pendulum and attractor fins. The anisotropic torque vanishes when the fins occupy 00, $50$ or $100\%$ of the available area and peaks when they they fill $25$ and $75\%$ of the available area. The peak at $75\%$ filling is slightly higher, presumably because of the increased masses of the pendulum and attractor fins. All of the numerical results presented in this section for the differential pendulum were obtained using fins that occupy $75\%$ of the available area. Figures 1, 2(b), and 3 all show differential pendulum geometries using the $75\%$ filling.
@@
Finally, we examined the dependence of the differential pendulum’s torque oscillations on the size of the $x\times x$ patch grid. This analysis was similar to that shown in Fig. 4(b) of Sec. III for the conventional geometry with $N$ pendulum and $N$ attractor fins. Here, we briefly summarize the key findings for the differential pendulum. Additional details are presented in the Supplemental Material.

### v3

@@ Abstract
Keywords: anisotropy, differential torsion pendulum, gravity, Yukawa
(Some figures may appear in colour only in the online journal)
Abstract
This Supplemental Material provides additional details about the numerical calculations used to characterize the proposed differential torsion pendulum. First, we present the geometry of pendulum and attractor fins and relate their positions to the calculated torque. Second, we discuss how anisotropy was incorporated into the short-range corrections to Newtonian gravity. Next, the dependence of the anisotropic torque on the angular size of the pendulum and attractor fins is presented. Finally, we comment on the accuracy of the numerically-calculated isotropic and anisotropic torques.
[…]
1 Introduction
@@ 2 Proposed Experiment
Figure 2(b), on the other hand, shows an attractor with $N$ fins and a pendulum with $N-1$ fins. In this case, to within a first order approximation, the net torque on the pendulum due to the isotropic component of the gravitational force vanishes for a large number of attractor orientations. As discussed below, the suppression of the isotropic component of the gravitational torque is enhanced as $N$ is increased. Note that the angular extent of the fins differ in Figs. 2(a) and (b). In (a), the fins occupy half of the available area and in (b) they occupy $75\%$ of the available area. The reason for using a larger fin size with the differential pendulum is discussed in Sec. 4.
@@ 3 Isotropic Gravity
The calculated torque has the expected periodicity of $2\pi/N$ and is consistent with an independent calculation of the net torque using the numerical methods described in Ref. [41].
@@ 4 Anisotropic Gravity
For the isotropic torque calculation presented in Sec. 3, the pendulum and attractor fins each occupied half of the total available area $\pi(R^{2}-r_{0}^{2})$. Figure 2(a) shows a top view of this geometry for the $N=4$ case. For the case of the differential torsion pendulum with $N-1$ pendulum fins and $N$ attractor fins, we found that the desired anisotropic torque vanished when the fins occupied half of the available area. The Supplemental Material presents a systematic study of the magnitude of the anisotropic torque as a function of the angular size of the pendulum and attractor fins. The anisotropic torque vanishes when the fins occupy 00, $50$ or $100\%$ of the available area and peaks when they they fill $25$ and $75\%$ of the available area. The peak at $75\%$ filling is slightly higher, presumably because of the increased masses of the pendulum and attractor fins. All of the numerical results presented in this section for the differential pendulum were obtained using fins that occupy $75\%$ of the available area. Figures 1, 2(b), and 3 all show differential pendulum geometries using the $75\%$ filling.
@@
Finally, we examined the dependence of the differential pendulum’s torque oscillations on the size of the $x\times x$ patch grid. This analysis was similar to that shown in Fig. 4(b) of Sec. 3 for the conventional geometry with $N$ pendulum and $N$ attractor fins. Here, we briefly summarize the key findings for the differential pendulum. Additional details are presented in the Supplemental Material.
@@ 6 Conclusion
S1 Differential Torsion Pendulum and Anisotropic Torque
Figure S1 shows a top view of a differential torison pendulum with $N-1=3$ pendulum fins and $N=4$ attractor fins with every fin subdivided into a grid of $x\times x$ patches, where $x=4$. The patch boundaries have been chosen such that all patches have the same area.
Figure S1: Top view of a pendulum with $N-1=3$ fins suspended above an attractor with $N=4$ fins. The vectors $\mathbf{r}_{\mathrm{p},i}^{\left(k\right)}$ and $\mathbf{r}_{\mathrm{a},j}^{\left(l\right)}$ locate the center-of-mass positions of patch $k$ in pendulum fin $i$ and patch $l$ in attractor fin $j$, respectively.
As shown in the figure, the fins have inner radii $r_{0}$ and outer radii $R$. For both the pendulum and attractor, the fins fill $75\%$ of the available area, corresponding to a filling fraction of $f=0.75$. In general, the angular widths $\beta_{\mathrm{p/a}}$ of the pendulum and attractor fins are given by
$\beta_{\mathrm{p}}$ $=\frac{2\pi f}{N-1}$ (S1)
$\beta_{\mathrm{a}}$ $=\frac{2\pi f}{N}.$ (S2)
Relative to the center of a fin, the angular position $\delta_{\mathrm{p/a}}^{\left(m\right)}$ of the center-of-mass of patch $m$ is given by
$\delta_{\mathrm{p/a}}^{\left(m\right)}=\frac{\beta_{\mathrm{p/a}}}{2}\left(\frac{2m+1}{x}-1\right),$ (S3)
where $m=0,1,\dots,x-1$ is an integer.
The radial boundaries of the patches are given by
$b_{n}=\sqrt{\left(\frac{n}{x}\right)R^{2}+\left(1-\frac{n}{x}\right)r_{0}^{2}},$ (S4)
where $n=0,1,\dots,x$ is another integer. The center-of-mass positions of patch $m$ along the radial direction, measured from the rotation axis, is then given by
$r^{\left(m\right)}=\frac{2}{3}\left(\frac{b_{m+1}^{3}-b_{m}^{3}}{b_{m+1}^{2}-b_{m}^{2}}\right).$ (S5)
Finally, the centre-of-mass position of patch $k$ in pendulum fin $i$ is given by
$\frac{\mathbf{r}_{\mathrm{p},i}^{\left(k\right)}}{r^{\left(k\right)}}=\cos\left(\theta+\delta_{\mathrm{p}}^{\left(k\right)}+\frac{2\pi i}{N-1}\right)\,\hat{\imath}+\sin\left(\theta+\delta_{\mathrm{p}}^{\left(k\right)}+\frac{2\pi i}{N-1}\right)\,\hat{\jmath}$ (S6)
and the centre-of-mass position of patch $l$ in attractor fin $j$ is given by
$\frac{\mathbf{r}_{\mathrm{a},j}}{r^{\left(l\right)}}=\cos\left(\phi+\delta_{\mathrm{a}}^{\left(l\right)}+\frac{2\pi j}{N}\right)\,\hat{\imath}+\sin\left(\phi+\delta_{\mathrm{a}}^{\left(l\right)}+\frac{2\pi j}{N}\right)\,\hat{\jmath}-z\,\hat{k}.$ (S7)
Equations (S6) and (S7) assume that the centre of mass of each pendulum fin lies in the $xy$-plane ($z=0$) and the centre of mass positions of the attractor fins are a distance $z$ below the $xy$-plane. If the pendulum and attractor fins each have a thickness $t$ and the top of the attractor fins are a distance $d$ below the bottom of the pendulum fins, then $z=-(t+d)$. The angles $\theta$ and $\phi$ give the orientations of the pendulum and attractor, respectively, relative to a reference $x$-axis. As discussed in the main manuscript, the force of patch $l$ in attractor fin $j$ on patch $k$ in pendulum fin $i$ due to Newtonian gravity is given by
$\mathbf{F}_{ij}^{\left(kl\right)}=\frac{Gm_{\mathrm{p}}m_{\mathrm{a}}}{\left(r_{ij}^{\left(kl\right)}\right)^{2}}\,\hat{r}_{ij}^{\left(kl\right)},$ (S8)
where $m_{\mathrm{p}}$ and $m_{\mathrm{a}}$ are the masses of the pendulum and attractor patches, respectively, $\mathbf{r}_{ij}^{\left(kl\right)}=\mathbf{r}_{\mathrm{a},j}^{\left(l\right)}-\mathbf{r}_{\mathrm{p},i}^{\left(k\right)}$, and $\hat{r}_{ij}^{\left(kl\right)}$ is a unit vector in the direction of $\mathbf{r}_{ij}^{\left(kl\right)}$.
The isotropic form of the Yukawa-corrected gravitational potential is given by [1]
$V_{ij}^{\left(kl\right)}=-\frac{Gm_{\mathrm{p}}m_{\mathrm{a}}}{r_{ij}^{\left(kl\right)}}\left(1+\alpha e^{-r_{ij}^{\left(kl\right)}/\lambda}\right).$ (S9)
Here, $\alpha$ is the strength correction parameter and $\lambda$ is the scale correction parameter. The isotropic form of the Yukawa-corrected gravitational force is then given by
$\mathbf{F}_{\mathrm{Y},ij}^{\left(kl\right)}=\frac{Gm_{\mathrm{p}}m_{\mathrm{a}}}{\left(r_{ij}^{\left(kl\right)}\right)^{2}}\Bigg{[}1+\alpha\left(1+\frac{r_{ij}^{\left(kl\right)}}{\lambda}\right)e^{-r_{ij}^{\left(kl\right)}/\lambda}\Bigg{]}\hat{r}_{ij}^{\left(kl\right)}.$ (S10)
We only consider gravitational anisotropies applied to the Yukawa term. The reason is that any correction (for example, from some theory of quantum gravity) would be a higher-order correction to the Einstein–Hilbert action. As the Einstein–Hilbert action preserves isotropy, we expect only the correction terms to break the isotropy of space. To make the Yukawa term anisotropic, we construct a normalized anisotropic unit vector given by
$\hat{r}^{\mathrm{a},\left(kl\right)}_{ij}(\varepsilon)=\frac{(1+\varepsilon)\sin\theta^{\prime\left(kl\right)}_{ij}\cos\phi^{\prime\left(kl\right)}_{ij}\,\hat{\imath}\,+\sin\theta^{\prime\left(kl\right)}_{ij}\sin\phi^{\prime\left(kl\right)}_{ij}\,\hat{\jmath}+\cos\theta^{\prime\left(kl\right)}_{ij}\,\hat{k}}{\sqrt{1+\varepsilon(2+\varepsilon)\sin^{2}\theta^{\prime\left(kl\right)}_{ij}\cos^{2}\phi^{\prime\left(kl\right)}_{ij}}}.$ (S11)
The angles $\theta^{\prime\left(kl\right)}_{ij}$ and $\phi^{\prime\left(kl\right)}_{ij}$ specify the direction of $\hat{r}^{\mathrm{a},\left(kl\right)}_{ij}(\varepsilon)$ in spherical coordinates and the parameter $\varepsilon$ determines the strength of the anisotropy. Note that, $\hat{r}^{\mathrm{a},\left(kl\right)}_{ij}\equiv\hat{r}^{\left(kl\right)}_{ij}$ when $\varepsilon=0$. The anisotropic Yukawa-corrected force can then be expressed as
$\mathbf{F}^{\mathrm{a},\left(kl\right)}_{\mathrm{Y},ij}=\frac{Gm_{\mathrm{p}}m_{\mathrm{a}}}{\left(r_{ij}^{\left(kl\right)}\right)^{2}}\Bigg{[}\hat{r}^{\left(kl\right)}_{ij}+\alpha\left(1+\frac{r_{ij}^{\left(kl\right)}}{\lambda}\right)e^{-r_{ij}^{\left(kl\right)}/\lambda}\,\hat{r}^{\mathrm{a},\left(kl\right)}_{ij}(\varepsilon)\Bigg{]}.$ (S12)
In a similar way, an anisotropic power-law corrected force can be expressed as
$\mathbf{F}^{\mathrm{a},\left(kl\right)}_{\mathrm{P},ij}=\frac{Gm_{\mathrm{p}}m_{\mathrm{a}}}{\left(r_{ij}^{\left(kl\right)}\right)^{2}}\left[\hat{r}_{ij}^{\left(kl\right)}+\left(\frac{k}{r_{ij}^{\left(kl\right)}}\right)\,\hat{r}^{\mathrm{a},\left(kl\right)}_{ij}(\varepsilon)\right].$ (S13)
Equations (S12) and (S13) reproduce Eqs. (2) and (3) of the main manuscript, respectively.
S2 Anisotropic Torque Amplitude Versus Filling Fraction
All of the results presented in Secs. S2 and S3 assume an anisotropic Yukawa-corrected gravitational force with $\alpha=1\times 10^{-3}$, $\lambda=1\rm\ mm$, and $\varepsilon=1\times 10^{-4}$. Furthermore, the pendulum and attractor fins were assumed to be made of tungsten ($\rho=19.3\rm\ g/cm^{3}$) and to have a thickness of $t=1\rm\ mm$ and inner and outer radii of $r_{0}=5\rm\ mm$ and $R=50\rm\ mm$, respectively. The distance separating the bottom of the pendulum and the top of the attractor was set to $d=50\rm\ \mu m$. All of these parameter values are consistent with those used in the main manuscript.
The first differential torsion pendulum geometry we considered used a fin filling fraction $f=0.5$ and reveled a vanishingly-small amplitude for the anisotropic component of the torque oscillation. This observation led us to investigate the dependence of the anisotropic amplitude on the filling fraction of the fins for $0\leq f\leq 1$. Figure S2 shows the results for differential pendulums with $N=20$, $30$, and $70$.
As expected, the anisotropic torque vanishes for $f=0$ (no pendulum and attractor fins) and $f=1$ (no empty space between neighboring fins). There is also a zero in the anisotropic torque amplitude at $f=0.5$, consistent with our original observation. The oscillation amplitude peaks between these minima at $f=0.25$ and $0.75$. In order to compare the results from the different values of $N$, the data in the figure were scaled to make the peak amplitudes at $f=0.75$ equal to one. When scaled in this way, Fig. S2 shows that oscillation amplitude’s dependence on $f$ is robust against changes to $N$.
Figure S2: Amplitude of the anisotropic component of the torque oscillation as a function of the pendulum and attractor fin filling fraction $f$. To facilitate easy comparisons, the data were scaled to make the peak height at $f=0.75$ equal to one. Data are shown for $N=20$, $30$, and $70$.
Notice that the peak at $0.75$ filling is slightly higher than that at $f=0.25$. One possible explanation is that the pendulum and attractor fins have higher total mass at the larger filling fraction which leads to greater torque amplitudes. Another possibility is that the difference in peak heights is due to changes in the patch mass as the fin size or, equivalently, the filling fraction is varied. For a given value of $N$, all of the points in Fig. S2 were calculated using the same $x\times x$ patch grid. For example, the $N=70$ data were calculated using $x=10$. Since the angular widths of the pendulum and attractor fins are proportional to $f$, the fins at $f=0.75$ are three times the size of the fins $f=0.25$. As a result, for fixed $x$, the patch masses are smaller for $f=0.25$ than they are for $f=0.75$. Intuitively, the smaller patch size is expected to lead to more accurate numerical results. As will be shown in Sec. S3, the amplitude of the torque oscillations decreases as the accuracy of the numerical calculations is increased. Therefore, the difference in peak heights at $f=0.25$ and $0.75$ could be due to the increased numerical accuracy of the calculations at lower filling fractions.
S3 Isotropic and Anisotropic Torque Amplitude Versus Number of Patches
This section is used to investigate the numerical accuracy of the torque calculations as a function of $x$ for the differential pendulum. The calculations where done using the parameters given in Sec. S2 and a filling fraction of $f=0.75$. The analysis is similar to that presented in the main manuscript for the conventional torsion pendulum with $N$ pendulum fins and $N$ attractor fins.
Figure S3 shows the amplitudes of the isotropic and anisotropic torque oscillations as a function of $x$. At the maximum value of $x$ used, the pendulum patch masses were less than or equal to $1.25\rm\ mg$. The attractor fins were divided into $x\times x$ patch grids using the same $x$ selected for the pendulum fins. Figure S3(a) shows the isotropic torque amplitude versus $x$ for a differential pendulum with $N=15$. Even at $x=81$ ($6561$ patches), the amplitude still shows an appreciable decrease as $x$ increases. Increasing $x$ from $57$ to $81$ (a doubling of the patch mass), causes the oscillation amplitude to decrease by $6.5\%$. The horizontal dashed line in the figure is a guide to the eye intended to highlight that the amplitude has not yet fully reached a stable value. The approach to stability slows even further as $N$ is increased. In Fig. 6 of the main manuscript, $N=34$ is the maximum value for which the isotropic oscillation amplitude is shown. Its percent decrease when the target pendulum patch mass is decreased from $2.5$ to $1.25\rm\ mg$ was $63\%$. Because of the prohibitive computation time, we did not attempt to reduce the target patch mass any further.
In contrast, the anisotropic oscillation amplitudes approached stability more quickly, even when considering large values of $N$. Figure S3(b) shows the $N=50$ case using the same target patch masses used in part (a) of the figure. As shown by the dashed line, there is minimal change ($2\%$) in the calculated amplitude as $x$ is increased from $31$ to $43$. Even when $N$ is increased to $100$, the change in the anisotropic oscillation amplitude is less then $7\%$ when the patch mass is reduced from $2.5$ to $1.25\rm\ mg$.
(a)Refer to caption
(b)Refer to caption
Figure S3: (a) Amplitude of the isotropic component of the torque oscillation as a function of $x$ for $N=15$. (b) Amplitude of the anisotropic component of the torque oscillation as a function of $x$ for $N=50$. The dashed horizontal lines are guides to the eye.